In the final stages of a moon landing, a lunar module descends under retro-thrust of its descent engine to a height of 𝑠 = 4 m above the lunar surface where it has a downward velocity of 4 m/s. Calculate the impact velocity with the moon’s surface if: The acceleration under the combined effect of gravity and retro-thrust is the following function of height 𝑎 = 𝑠/2.


1
Expert's answer
2022-05-07T14:40:40-0400

"v_2=\sqrt{v_1^2+2as}=5.7~\frac ms."